OpenStudy (anonymous):

Simplify the number using the imaginary unit i. radical -144 A. 12 B. -12 C. 12i D. 144i

OpenStudy (anonymous):

looks like this\[\sqrt{-144}\]

OpenStudy (anonymous):

D

OpenStudy (anonymous):

\[i=\sqrt{-1}\]

OpenStudy (anonymous):

oh wait that makes it C actually

OpenStudy (anonymous):

\[\Large \sqrt{-1*12^2} = 12\sqrt{-1}=12i\]

OpenStudy (anonymous):

Thank you very much.